A merchant sells one type of tea for $5.50 per pound and a second type for $4.25 per pound. She wishes to mix the tea to form a 60 lb mixture that will sell for $4.55 per pound. How many pounds of each type of tea should be used in the mixture?
Answer by stanbon(75887)

A merchant sells one type of tea for $5.50 per pound and a second type for $4.25 per pound. She wishes to mix the tea to form a 60 lb mixture that will sell for $4.55 per pound. How many pounds of each type of tea should be used in the mixture?
--------------------
Quantity:: a + b = 60 lbs
Value::: 550a+425b = 455*60
-------------------------------
Modify for elimination:
550a + 550b = 550*60
550a + 425b = 455*60
---------
Subtract and solve for "b":
125b = 95*60
b = (19/25)60
b = 45.6 lbs (amt. of $4.25 tea needed)
------------
a = 60-45.6 = 14.4 lbs (amt of $5.50 tea needed)
===============
Cheers,
Stan H.
------------